Circular No. 6212

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1735–27 B. Aschenbach, W. Voges, and W. Becker, Max-Planck-Institut für

Extraterrestrische Physik, Garching, report: “Inside the GRO J1735–27 location error box (IAUC 6207), we have found a bright x-ray source, RX J1735.9–2726, in the ROSAT all-sky survey data centered on α = 17^{h}35^{m}59^{s}.0, δ = −27^{o}26′22′′ (equinox 2000.0; estimated error 20′′, 1σ radius). The PSPC count rate in the ROSAT band (0.1–2.4 keV) was 29 counts/s during 1990 Sept. 5–7. In ROSAT observations between 1992 Feb. 29 and Mar. 3, RX J1735.9–2726 yielded 0.0031 PSPC counts/s; during 1992 Sept. 12–16, the source was below 0.0025 HRI counts/s (1σ upper limit). The 1990 energy spectrum is consistent with a power law with photon spectral index −2.0 ± 0.4 and a normalization of 1.1 photon cm^{−2} s^{−1} keV^{−1} at 1 keV. The N_{H} absorption column density is (10.5 ± 1.5) × 10^{21} cm^{−2}. The absorption-corrected energy flux is 5.6 × 10^{−9} erg cm^{−2} s^{−1}. A search for pulsations in a range bracketing the BATSE period (IAUC 6207) revealed only weak evidence for a periodic signal at P = 4.4377 s (mean epoch JD 2448140.707024 TDB). RX J1735.9–2726 clearly is a transient with a hard spectrum consistent with that reported for GRO J1735–27 (and similar brightness), suggesting that the two sources are identical. Within a radius of 1′ around RX J1735.9–2726, the SIMBAD database contains only IRAS 17328–2725. We note that RX J1735.9–2726 is offset by 16′.2 from (but only slightly outside of the error box of) the bright, time-variable, x-ray source GS 1734–275 detected with the Ginga satellite in 1987 and observed until 1990 (IAUC 4571; Yamauchi and Koyama 1990, PASJ 42, L83). The energy spectrum of GS 1734–275 is well matched by a blackbody spectrum of kT = 0.58 keV. A blackbody spectral fit to RX J1735.9–2726 is acceptable with kT = (0.5 ± 0.1) keV.”

SUPERNOVA 1995X IN UGC 12160 P. Sicoli reports the following precise position for SN 1995X (cf. IAUC

6211) obtained at Sormano, Italy, by M. Cavagna and F. Manca on Aug. 24.863 UT with the 0.5-m reflector (+ CCD): α = 22^{h}40^{m}51^{s}.30, δ = +75^{o}10′11′′.5 (equinox 2000.0; GSC reference stars).

COMET C/1995 Q1 (BRADFIELD) Visual m_{1} estimates: Aug. 21.36 UT, 5.5 (P. Camilleri, Cobram, Vic.,

20×80 binoculars); 25.39, 5.5 (P. Caldwell, Geelong, Vic., 7×50 binoculars).

1995 August 25 Daniel W. E. Green